An asset costs 14,000 dollars and has a scrap value of 2,400 dollars after seven years. Calculate its book value using straight-line depreciation (Perform all calculations using 5 significant figures and round your answer to the nearest dollar).

(a) After one year?

(b) After four years?

(c) After seven years?
Answer #1

Answer

Depreciation rate per year =( Asset cost -scrap value)/life

=(14000-2400)/7

=1657.14286

Book value after n years =asset cost - n * depreciation rate per year

========

(a) After one year?

=14000-1657.14286=12342.8571=12343

(b) After four years?

=14000-4*1657.14286=7371.42856=7371

(c) After seven years?

=14000-7*1657.14286=2399.99998=2400